# “It’s Always a Losing Game”: How Workers Understand and Resist Surveillance Technologies on the Job

### v1

@@ 4.2.2. Research
Doing research was another accessible technique that workers employed. Some workers tried to increase their awareness of workplace surveillance by reading news articles and doing online searches about the topic. Workers on Reddit regularly posted news articles about workplace surveillance, such as to share the latest technologies in the space [R6], as well as companies that were actively surveilling their workers [R105, 192]. P1 reported reading news articles about how companies surveil workers without disclosing it and now assumed that most companies practice it. P3 used YouTube and other sources to learn more about Hubstaff and its capabilities. Through his research, P3 learned that his supervisor had the ability to take screenshots of his screen at any time without notifying him. Both P3 and P5 also tried to look into research studies about workplace surveillance but P5 acknowledged that she found little information available, especially regarding what her rights were as a worker:
@@ 4.2.3. Obfuscation and Disablement
Workers also adopted strategies to keep their work and personal files separate, such as through virtual machines or separate devices in order to keep their personal information and activities outside of the purview of management. P3, for example, stated that he kept his social media and banking apps on his mobile phone. While some workers had access to multiple computers that allowed them to keep work and personal files separate, P5, who was not given a separate work computer by her company, had to resort to using her personal computer. One way she kept personal and work files separate was to create a separate user account exclusively for work-related purposes on her personal computer. However, P5 acknowledged that sometimes it is difficult to separate the two: “Sometimes I do find that some things collide, just because, while I’m working, I have to do personal tasks. But that has helped me a lot in just mentally separating those things, and not mixing my work so much with my life-related duties.”
@@
Workers also disabled technologies while on the job. For example, P1 would sometimes refuse to turn her camera on, especially on days when she was not feeling well. However, she was careful not to make it an everyday occurrence, since she would be at risk of being fired and saw other coworkers be fired for refusal: “I have told them that my camera wasn’t working on multiple occasions […] I’ve shut that little camera door and told them my camera wasn’t working. I’ve told them that I lost my webcam when I moved.” However, P1 noted that certain surveillance technologies like keyloggers were harder to bypass and many workers did not have the ability to disable monitoring systems. P5 recounted the moment she realized her tracker could not be disabled, stating, “That’s when I realized, Oh, this is not optional. I just have to deal with this. And I have to suck it up if I want to get paid.” Although some workers were able to successfully avoid, trick, and even stop the surveillance they were subjected to without detection, in a heavily surveilled workplace, there was often little room for such maneuvers, forcing workers to explore alternative strategies.
@@ 4.2.4. Ignoring and Soldiering
P9 caught on that if she got her work done earlier, that they would give her more work to do, so she intentionally did not work at 100% to not set a new productivity norm:
@@
R134, the same manufacturing worker who previously stated that they experienced anxiety whenever their monitoring software showed a red screen eventually learned to ignore it, adding: “Even though I can’t turn the screen off, I learned to turn the screen to the back of the machine and work at my own pace. I know this is purely manipulation, they will not fire us because we are skilled operators that work for small wages and produce lots of profits.” Knowing that their employer may have difficulty replacing them due to their expertise gave this worker a sense of power and agency to control their pace of work in the face of constant surveillance.
@@ 4.2.5. Legal Action and Collective Organizing
Although some posters on Reddit suggested resisting through unionizing or taking legal action, many of them admitted the difficulties in doing so. In cases where workers were protected through a union, some found that it did little to change their conditions when it came to monitoring. P2, who was represented by a union noted that oftentimes, the union would not result in any substantial change, especially when workers are on short-term contracts:
@@
Some workers on Reddit described facing strong anti-union sentiment among their coworkers, while others mentioned the trouble of living in a “right-to-work” state. As R134 put it, “if you mention unionizing, you are immediately fired.” They added that although workers are legally protected against retaliation for unionizing, their employers “usually come up with some excuse about attendance or inefficiency and then just like that, you’re out.” Others noted the difficulties of taking legal action for employment issues. R147 added that they were “surprised by how many people suggest to just hire a lawyer and sue.” They described the barriers their spouse experienced while attempting to seek legal counsel in a wage theft case. Because the company used a time tracking system, the poster assumed that it would be an clear-cut case. However, of the twenty employment attorneys R147 reached out to, most did not respond and the few who did told them that it wasn’t worth the time and effort: “My partner and I are both well-educated and earn reasonable incomes, yet we’ve found it extremely challenging to access the legal system. I can only imagine how much harder it must be for others. In my experience, the legal system provides minimal protections for workers and largely serves to reinforce the oppressive system we live in.” These testimonials illustrate the difficulties of collective organizing against workplace surveillance, underscoring the lack of meaningful worker protections against its use and the need for alternative approaches.
@@ 4.2.6. Quitting
Ultimately, in cases where workers felt like they had no alternatives, many indeed chose to quit their jobs to get away from intrusive monitoring. Quitting was the most frequently named resistance tactic suggested by posters on Reddit to deal with workplace surveillance. During our interviews, P1 noted the limited amount of power workers have to combat surveillance where they are usually faced with only two choices: “You can either quit, or you can stay there and try to get a paycheck and be uncomfortable.” However, workers stated how their awareness or experiences with workplace surveillance informed how they performed their job search, explicitly choosing to work in jobs with less invasive surveillance and more worker autonomy and refusing to work in jobs that did not have this in place. Together, these resistance tactics, including quitting, illustrate how workers showed not only their autonomy, but also their creativity and ingenuity in the face of constant surveillance and gain back a little control over their working conditions.
@@ 4.3.1. Stronger Governmental Protections
P1, a US-based worker, looked to Norway as a model for how workplace surveillance should be regulated, believing that the country has implemented stronger data privacy protections as compared to the US. In addition to enforcing the EU’s GDPR, Norway introduced the Working Environment Act which limits how employers can access their employees emails and camera surveillance (IAPP, 2022). Other workers called for stronger regulation that would place hard limits on monitoring, provide more transparency around its use, and allow workers to provide meaningful consent and the ability to opt-out of surveillance. More specifically, P1 and P3 wanted to limit the use of more excessive forms of monitoring, such as keyboard, browser, and mouse tracking, as this type of data can be easily misused for purposes other than tracking one’s productivity. P3 also called for limitations on how personal data is collected and stored by these technologies, advocating for data to be deleted within a year or after a project is over. Additionally, P3 wanted institutions to have clearer policies around handling worker harassment as a result of surveillance: “They can misuse this in a very bad way where they can micromanage […] They can actually see exactly what [you’re] working on every second […] They can just use it [for] harassment. So there should be some way of telling your manager that you should not be doing so and so things which are detrimental to an employee, that [prevents] harassment.”
@@
However, some workers expressed more skeptical views on governmental regulation and pointed to its limitations. P5, who worked in Mexico for a US-based company as a remote worker believed that the US did not have to adhere to labor laws specific to Mexico, which she claimed contributed to poorer working conditions and no paid time off:“The current condition that I’m in, in my current employment, and the way that they operate should be illegal […] What measures are they taking in order to comply to the legal standards of the countries that they are hiring at?” P8, who was based in the Philippines but served clients in the US was also skeptical of stronger regulation due to distrust of her own government since, she believed, the Philippines did not have a great track record of protecting their citizen’s privacy and security: “[In the] Philippines […] we don’t have a solid foundation regarding security […] If you’re asking if I should have [regulation], for now, no, since we don’t have a solid foundation regarding that security.” P6, a US-based worker, also did not believe that regulation would stop companies because of the belief that they worked in direct collaboration with politicians and would continue to do what they wanted, regardless of regulation: “They’re never gonna not collect a certain amount of information on you, because I feel a certain way. So no, they’re these corporations will have free rein to do essentially, whatever they want.” While many workers called for stronger government intervention, others were skeptical that regulation alone could adequately curb invasive monitoring practices, especially as workplaces become more fissured and global.
@@ 4.3.2. Stronger Organizational Policies
Beyond regulation, interviewees wanted their workplaces to foster a culture of trust, with an understanding that workers do not have to perform at 100% at all times and also have the ability to dictate the conditions in which they work best. R151 said that it was actually poor management practices that makes these tracking systems so popular: “Easily measurable factors aren’t necessarily good indicators of employee performance, while meaningful indicators are often difficult to quantify. In my view, these tracking systems stem from misguided, careless, and lazy management.” P2 added that workers would be more driven to be productive “if they feel like they’re trusted, and in a place that they’re happy to work…if you’re being watched at all times, it’s not the best environment.” P3, who was also a manager in his full-time role, stated that he fully trusted his employees, even during periods where they may not be as productive:
@@
Building this culture of trust would lessen the general need for WSTs, especially those that continuously track workers by the minute. However, workers and managers acknowledged that some forms of monitoring can be helpful for training and accountability purposes. In such cases, workers wanted to have input on how surveillance is implemented and to determine what metrics are valuable to support the health of the business and workers. P6 stated that although companies do have a responsibility to improve their business, “employees should have a say in what the metrics are.” As part of ways to use these technologies to support their well-being, workers wanted these technologies to be more accepting of accommodations. P5 felt that workers should not be expected to perform 100% of the time and it should be up to the worker to determine if they want to use time-tracking software or not: “In normal jobs, you’re not expected to be productive 100% of the time. So why remote workers should feel compelled to do that or feel forced to do that, with this kind of tracking software?” Similarly, P3 did not believe that maintaining an 85% productivity score daily is sustainable in the long term. Instead, employees should be evaluated by their productive averages over a longer period of time:
There should be some slack of like, if I meet my productivity goals for one week, then I’ll have at least two chances for the next week. Or there are chances when accidentally we will come get below 85% They may be some sudden emergency or I forget to turn off things. So there should be some way of making the employee feel comfortable about his continuity in the project.” — P3
P6, who expressed more positive views on monitoring technologies said that it was because his company did a good job of supporting people who may fall behind on metrics. This included an extensive training period where metrics were not held against new workers. He also stated that workers were still able to manage their own schedules: “It most definitely starts from the top down, and so seeing that in practice has been a nice reassurance that they’re here to work with me, instead of just get rid of me, if I don’t have metrics.” Yet, P2 questioned the possibility that WSTs could be built to help to support workers. They believed that these technologies can never be truly ethical unless the issue of treating humans like machines is addressed first:
@@ 5. Discussion
Our findings reveal how WSTs created a culture of distrust and anxiety that negatively impact both workers and managers. Workers reported fears of privacy invasion and disciplinary action, exacerbated by opaque and unaccommodating monitoring practices. Managers were also subjected to surveillance and experienced paranoia as a result of WSTs. Despite having limited power and legal protections, workers resisted WSTs through tactics like commiseration, obfuscation, and ”soldiering,” which emphasized their refusal over compliance. Collective action, though challenging, emerged subtly through the sharing of strategies and growing political consciousness on Reddit. By drawing inspiration from these tactics, we call on the CSCW community to take an emancipatory approach to worker-centered design that supports worker autonomy, resistance, and collective action.
@@ 5.1. The Losing Game of a Surveilled Workplace
Throughout our interviews and in Reddit posts, workers described the multiple fears and anxieties they experienced as a result of being subjected to WSTs, even ones perceived to be less ”intrusive” or ”harmful” such as productivity tracking (Constantinides and Quercia, 2022). For remote workers, the ability for companies to surveil them in their own homes created unique privacy concerns not seen in other forms of work involving an overt and often unintended spillage of their private lives into the workplace. Workers were not only concerned about the extensive personal data being collected about them, but also that these tools were primarily used for discipline and punishment rather than support. P1 and P2’s experiences with Emotion AI systems demonstrate how workers were placed under a punitive three-strikes-and-you’re-fired policy any time the system flagged them for not adhering to inflexible and unrealistic techniques of the body (Mauss, 1973). As P1 noted, this created a “big snowball effect” that made workers more susceptible to additional surveillance, suspicion, and harassment by management.
@@
Together, these examples show how WSTs created a reverberating culture of distrust that hurt workers and managers alike. The punitive nature of WSTs reinforced divisions within the workplace, where supervisors assumed the worst in their employees, workers felt resentment towards their bosses, and coworkers were pitted against each other. At the root of this dynamic is the dehumanization that occurs when workers and their labor are abstracted into measurable units of accountability and compared against the metrics of others (Haggerty and Ericson, 2000). This in turn disincentivizes workers and bosses alike from investing in the harder-to-measure interpersonal aspects of work, like building trust and open communication that are crucial to creating a healthy work environment(Thornburg, 2015). Similar to Veena Dubal’s description of laboring under Uber’s algorithmic management as “gambling” when “the house always wins” (Dubal, 2023), we echo the sentiment shared by the Reddit poster who introduced this paper — that WSTs pull workers and even their managers into “a losing game” where nobody wins.
According to Gould (Gould, 2024; Gould et al., 2023), addressing these challenges as HCI researchers requires critical examination of the epistemologies of work that underpin measurement frameworks of WSTs, as well as the assumptions that guide our understandings and critiques of these systems. We call for a shift toward worker-centered design methodologies (Fox et al., 2020), where workers are actively involved in envisioning and designing tools and frameworks that can genuinely reflect and support the complexity and variability inherent in their work. Only by prioritizing worker voices over ”stochastic machine witnesses” (Gould, 2024) — automated systems that reduce human activity to decontextualized metrics — can HCI researchers design systems that could truly enhance the workplace experience where both workers and organizations might thrive.

### v2

@@ 4.2.2. Research
Doing research was another accessible technique that workers employed. Some workers tried to increase their awareness of workplace surveillance by reading news articles and doing online searches about the topic. Workers on Reddit regularly posted news articles about workplace surveillance, such as to share the latest technologies in the space [R6], as well as companies that were actively surveilling their workers [R105, 192]. P1 reported reading news articles about how companies surveil workers without disclosing it and now assumed that most companies practice it. P3 used YouTube and other sources to learn more about Hubstaff and its capabilities. Through his research, P3 learned that his supervisor had the ability to take screenshots of his screen at any time without notifying him. Both P3 and P5 also tried to look into studies about workplace surveillance, but P5 acknowledged that she found little information available, especially regarding what her rights were as a worker:
@@ 4.2.3. Obfuscation and Disablement
Remote workers also adopted strategies to keep their work and personal files separate, such as through virtual machines or separate devices in order to keep their personal information and activities outside the purview of management. P3, for example, stated that he kept his social media and banking apps on his mobile phone. While some workers had access to multiple computers that allowed them to keep work and personal files separate, P5, who was not given a separate work computer by her company, had to resort to using her personal computer. One way she kept personal and work files separate was to create a separate user account exclusively for work-related purposes on her personal computer. However, P5 acknowledged that sometimes it is difficult to separate the two: “Sometimes I do find that some things collide, just because, while I’m working, I have to do personal tasks. But that has helped me a lot in just mentally separating those things, and not mixing my work so much with my life-related duties.”
@@
Workers also disabled technologies while on the job. For example, P1 would sometimes refuse to turn her camera on, especially on days when she was not feeling well. However, she was careful not to make it an everyday occurrence, since she would be at risk of termination and saw coworkers fired for refusing to enable theirs: “I have told them that my camera wasn’t working on multiple occasions […] I’ve shut that little camera door and told them my camera wasn’t working. I’ve told them that I lost my webcam when I moved.” However, P1 noted that certain surveillance technologies like keyloggers were harder to bypass and many workers did not have the ability to disable monitoring systems. P5 recounted the moment she realized her tracker could not be disabled, stating, “That’s when I realized, Oh, this is not optional. I just have to deal with this. And I have to suck it up if I want to get paid.” Although some workers were able to successfully avoid, trick, and even stop the surveillance they were subjected to without detection, in a heavily surveilled workplace, there was often little room for such maneuvers, forcing workers to explore alternative strategies.
@@ 4.2.4. Ignoring and Soldiering
P9 caught on that if she got her work done earlier, that they would give her more work to do, so she intentionally slowed her pace to avoid setting a new productivity norm:
@@
R134, the same manufacturing worker who previously stated that they experienced anxiety when their monitoring software showed a red screen eventually learned to ignore it, adding: “Even though I can’t turn the screen off, I learned to turn the screen to the back of the machine and work at my own pace. I know this is purely manipulation, they will not fire us because we are skilled operators that work for small wages and produce lots of profits.” Knowing that their employer may have difficulty replacing them due to their expertise gave this worker a sense of power and agency to control their pace of work in the face of constant surveillance.
@@ 4.2.5. Legal Action and Collective Organizing
Although some posters on Reddit suggested resisting through unionization or taking legal action, many of them admitted the difficulties in doing so. In cases where workers were protected through a union, some found that it did little to change their conditions when it came to monitoring. P2, who was represented by a union, noted that bringing this issue to their representatives had not led to substantial change, especially for workers on short-term contracts:
@@
Some workers on Reddit described facing strong anti-union sentiment among their coworkers, while others mentioned the trouble of living in a “right-to-work” state. As R134 put it, “if you mention unionizing, you are immediately fired.” They added that although workers are legally protected against retaliation for unionizing, their employers “usually come up with some excuse about attendance or inefficiency and then just like that, you’re out.” Others noted the difficulties of taking legal action for employment issues. R147 added that they were “surprised by how many people suggest to just hire a lawyer and sue.” They described the barriers their spouse experienced while attempting to seek legal counsel in a wage theft case. Because the company used a time tracking system, the poster assumed that it would be a clear-cut case. However, of the twenty employment attorneys R147 reached out to, most did not respond and the few who did told them that it wasn’t worth the time and effort: “My partner and I are both well-educated and earn reasonable incomes, yet we’ve found it extremely challenging to access the legal system. I can only imagine how much harder it must be for others. In my experience, the legal system provides minimal protections for workers and largely serves to reinforce the oppressive system we live in.” These testimonials illustrate the difficulties of collective organizing against workplace surveillance, underscoring the lack of meaningful worker protections against its use and the need for alternative approaches.
@@ 4.2.6. Quitting
Ultimately, in cases where workers felt like they had no alternatives, many indeed chose to quit their jobs to get away from intrusive monitoring. Quitting was the most frequently named resistance tactic suggested by posters on Reddit to deal with workplace surveillance. During their interview, P1 noted the limited amount of power workers have to combat surveillance, faced with only two real choices: “You can either quit, or you can stay there and try to get a paycheck and be uncomfortable.” However, workers stated that their awareness of and experiences with surveillance in previous workplaces informed their subsequent job searches, explicitly choosing jobs with less invasive monitoring and more worker autonomy, and avoiding companies where this was not the case. Together, these resistance tactics, including quitting, illustrate workers’ ingenuity in seeking to gain control over their working condition, even in the face of constant surveillance.
@@ 4.3.1. Stronger Governmental Protections
P1, a US-based worker, looked to Norway as a model for how workplace surveillance should be regulated, believing that the country has implemented stronger data privacy protections as compared to the US. In addition to enforcing the EU’s GDPR, Norway introduced the Working Environment Act which limits how employers can access their employees emails and camera surveillance (IAPP, 2022). Other workers called for stronger regulation that would place hard limits on monitoring, provide more transparency around its use, and allow workers to provide meaningful consent and the ability to opt-out of surveillance. More specifically, P1 and P3 called for limits on more excessive forms of monitoring, such as keyboard, browser, and mouse tracking, as this type of data can be easily misused for purposes other than tracking one’s productivity. P3 also called for limitations on how personal data is collected and stored by these technologies, advocating for data to be deleted within a year or after a project is over. Additionally, P3 wanted institutions to have clearer policies around handling worker harassment as a result of surveillance: “They can misuse this in a very bad way where they can micromanage […] They can actually see exactly what [you’re] working on every second […] They can just use it [for] harassment. So there should be some way of telling your manager that you should not be doing so and so things which are detrimental to an employee, that [prevents] harassment.”
@@
However, some workers expressed more skeptical views on regulatory approaches and pointed to limitations. P5, who worked in Mexico for a US-based company as a remote worker believed that the US did not have to adhere to labor laws specific to Mexico, which she claimed contributed to poorer working conditions and no paid time off:“The current condition that I’m in, in my current employment, and the way that they operate should be illegal […] What measures are they taking in order to comply to the legal standards of the countries that they are hiring at?” P8, who was based in the Philippines but served clients in the US was also skeptical of stronger regulation due to distrust of her own government since, she believed, the Philippines did not have a great track record of protecting their citizen’s privacy and security: “[In the] Philippines […] we don’t have a solid foundation regarding security […] If you’re asking if I should have [regulation], for now, no, since we don’t have a solid foundation regarding that security.” P6, a US-based worker, also did not believe that regulation would stop companies because of the belief that they worked in direct collaboration with politicians and would continue to do what they wanted, regardless of regulation: “They’re never gonna not collect a certain amount of information on you, because I feel a certain way. So no, they’re these corporations will have free rein to do essentially, whatever they want.” While many workers called for stronger government intervention, others were skeptical that regulation alone could adequately curb invasive monitoring practices, especially as workplaces become more fissured and global.
@@ 4.3.2. Stronger Organizational Policies
Beyond regulation, interviewees wanted their workplaces to foster a culture of trust, with an understanding that workers do not have to perform at 100% at all times and also have the ability to dictate the conditions in which they work best. R151 said that it was actually poor management practices that make these tracking systems so popular: “Easily measurable factors aren’t necessarily good indicators of employee performance, while meaningful indicators are often difficult to quantify. In my view, these tracking systems stem from misguided, careless, and lazy management.” P2 added that workers would be more driven to be productive “if they feel like they’re trusted, and in a place that they’re happy to work…if you’re being watched at all times, it’s not the best environment.” P3, who was also a manager in his full-time role, stated that he fully trusted his employees, even during periods where they may not be as productive:
@@
To those we spoke with, building this culture of trust would lessen the general need for WSTs, especially those that continuously track workers by the minute. However, workers and managers acknowledged that some forms of monitoring can be helpful for training and accountability purposes. In such cases, workers wanted to have input on how surveillance is implemented and to determine what metrics are valuable to support the health of the business and workers. P6 stated that although companies do have a responsibility to improve their business, “employees should have a say in what the metrics are.” As part of ways to use these technologies to support their well-being, workers wanted these technologies to be more accepting of accommodations. P5 felt that it should be up to the worker to determine whether to use time-tracking software: “In normal jobs, you’re not expected to be productive 100% of the time. So why remote workers should feel compelled to do that or feel forced to do that, with this kind of tracking software?” Similarly, P3 did not believe that maintaining an 85% productivity score daily is sustainable in the long term. Instead, employees should be evaluated by their productive averages over a longer period of time:
“There should be some slack of like, if I meet my productivity goals for one week, then I’ll have at least two chances for the next week. Or there are chances when accidentally we will come get below 85%. They may be some sudden emergency or I forget to turn off things. So there should be some way of making the employee feel comfortable about his continuity in the project.” — P3
P6, who expressed more positive views on monitoring technologies, said that it was because his company did a good job of supporting people who may fall behind on metrics. This included an extensive training period where metrics were not held against new workers. He also stated that workers were still able to manage their own schedules: “It most definitely starts from the top down, and so seeing that in practice has been a nice reassurance that they’re here to work with me, instead of just get rid of me, if I don’t have metrics.” Yet, P2 questioned the possibility that WSTs could be built to support workers. They believed that these technologies can never be truly ethical unless the issue of treating humans like machines is addressed first:
@@ 5. Discussion
Our findings reveal how WSTs create a culture of distrust and anxiety that negatively impacts both workers and managers. Workers reported fears of privacy invasion and disciplinary action, exacerbated by opaque and unaccommodating monitoring practices. Managers were also subjected to surveillance and experienced paranoia as a result of WSTs. Despite having limited power and legal protections, workers resisted WSTs through tactics like commiseration, obfuscation, and “soldiering,” which emphasized their refusal over compliance. Collective action, though challenging, emerged subtly through the sharing of strategies and growing political consciousness on Reddit. By drawing inspiration from these tactics, we call on the CSCW community to take an emancipatory approach to worker-centered design that supports worker autonomy, resistance, and collective action.
@@ 5.1. The Losing Game of a Surveilled Workplace
Throughout our interviews and in Reddit posts, workers described the multiple fears and anxieties they experienced as a result of being subjected to WSTs, even ones perceived to be less “intrusive” or“”harmful” such as productivity tracking (Constantinides and Quercia, 2022). For remote workers, the ability for companies to surveil them in their own homes created unique privacy concerns not seen in other forms of work involving an overt and often unintended spillage of their private lives into the workplace. Workers were not only concerned about the extensive personal data being collected about them, but also that these tools were primarily used for discipline and punishment rather than support. P1 and P7’s experiences with emotion AI systems demonstrate how workers were placed under a punitive three-strikes-and-you’re-fired policy any time the system flagged them for not adhering to inflexible and unrealistic techniques of the body (Mauss, 1973). As P1 noted, this created a “big snowball effect” that made workers more susceptible to additional surveillance, suspicion, and harassment by management.
@@
Together, these examples show how WSTs created a reverberating culture of distrust that hurt workers and managers alike. The punitive nature of WSTs reinforced divisions within the workplace, where supervisors assumed the worst in their employees, workers felt resentment towards their bosses, and coworkers were pitted against each other. At the root of this dynamic is the dehumanization that occurs when workers and their labor are abstracted into measurable units of accountability and compared against the metrics of others (Haggerty and Ericson, 2000). This in turn disincentivizes workers and bosses alike from investing in the harder-to-measure interpersonal aspects of work, like building trust and open communication that are crucial to creating a healthy work environment(Nuñez, 2015). Similar to Veena Dubal’s description of laboring under Uber’s algorithmic management as “gambling” when “the house always wins” (Dubal, 2023), we echo the sentiment shared by the Reddit poster who introduced this paper — that WSTs pull workers and even their managers into “a losing game” where nobody wins.
According to Gould (Gould, 2024; Gould et al., 2023), addressing these challenges as HCI researchers requires critical examination of the measurement frameworks that underpin WSTs, as well as the assumptions that guide our understandings and critiques of these systems. We call for a shift toward worker-centered design methodologies (Fox et al., 2020), where workers are actively involved in envisioning and designing tools and frameworks that can genuinely reflect and support the complexity and variability inherent in their work. Only by prioritizing worker voice over “stochastic machine witnesses” (Gould, 2024) — automated systems that reduce human activity to decontextualized metrics — can HCI researchers design systems to enhance the workplace experience such that both workers and organizations might thrive.